OpenStudy (aarontookold):

A combined circuit has two resistors in parallel (10.0 ohms and 14.0 ohms) and another in series (5.0 ohms). If the power source is 9.0 volts, what is the current? I WILL FAN AND MEDAL

OpenStudy (scooby6363):

Find the equivalent resistance for the circuit by first finding the resistance of the parallel portion of the circuit. This will be: (1/10 + 1/14) = .1714 Now do the inverse of this number by doing 1/.1714 = 5.83 ohms. This will be the resistance of the parallel portion of the circuit. Now add the resistance of the parallel portion with the series resistance which is only 5 ohms. 5.83 + 5 = 10.83 ohms. Current = Voltage/Resistance I = 9.0 volts / 10.83 ohms = .831 amperes This will be the total current of the circuit.

OpenStudy (physics_dude99):

 The current will split and then recombine. The parallel combination will have an equivalent resistance less than the smaller of the resistances. The total power consumed by the circuit is P = I*V